V. Response of the Eye. The human eye is most sensitive to green light of wavelength 505 nm (505
x10-9 m). Experiments have found that when people are kept in the dark room until their eyes adapt to
the darkness, a single photon of green light will trigger receptor cells in the rods of the retina.
(a) What
is the frequency of this photon?
(b) How much energy (in joules and electron volts) does it deliver to
the receptor cells? (c) to appreciate what a small amount of energy this is, calculate how fast a typical
bacterium of mass 9.5x10-12 g would move if it had that much energy.

VI. A photon of yellow light have a wavelength of 560 nm. Find the photon’s frequency, magnitude of
momentum, and energy. Express the energy in both joules and electron volts. ( 1eV=1.602x10-19J)
1
Expert's answer
2020-06-08T10:29:46-0400

Frequency is calculate by this formula "\frac{c}{\lambda}=\frac{3\times10^8}{505\times10^{-9}}=5.94\times10^{14}Hz"


Energy ="hv=6.63\times10^{-34}\times5.94\times10^{14}=3.9\times10^{-19}J"